Five years ago, John's age was half of the age he will be in 8 years. How old is he now?
Answer by Theo(13342)

let x be john's age now.
five years ago, john's age was x-5
eight years from now, john's age will be x+8
-----
Five years ago, John's age was half of the age he will be in 8 years.
this means that (x-5) = (x+8)/2
-----
solving for x, we get:
2 * (x-5) = x + 8
which becomes:
2x - 10 = x + 8
which becomes:
x = 18
-----
john is 18 years old now.
5 years ago john was 13
8 years from now john will be 26
13 = 1/2 * 26
-----